OpenStudy (anonymous):

Suppose the probability that it rains in the next two days is one-third for tomorrow and one-sixth for the day after tomorrow. What is P(rain tomorrow

OpenStudy (anonymous):

assuming they are independent it is \[\frac{1}{3}+\frac{1}{6}-\frac{1}{3}\times \frac{1}{6}\]

OpenStudy (anonymous):

?